# Antenna Position Optimization for Movable Antenna-Empowered Near-Field Sensing

### v2

@@ I Introduction
In future sixth-generation (6G) wireless systems, significant advancement in their communication and sensing capabilities is envisioned[1, 2]. A multitude of research efforts have been dedicated to achieving ultra-high transmission rates and precise information acquisition from the physical environment, driven by the demands of emerging applications such as smart healthcare, vehicle-to-everything (V2X) and virtual reality (VR) [3, 4]. However, existing communication and/or sensing systems typically deploy fixed-position antennas (FPAs) at the transmitter/receiver (Tx/Rx), which cannot fully exploit the spatial degrees of freedom (DoFs) and may only achieve a sub-optimal performance.
@@
Notations: Boldface lower and upper cases are used to represent vectors and matrices, respectively. The conjugate of a complex number is denoted by $(\cdot)^{*}$. The transpose and conjugate transpose of a matrix/vector are denoted by $(\cdot)^{\top}$ and $(\cdot)^{\mathrm{H}}$, respectively. The sets of $N_{1}\times N_{2}$ dimensional real and complex matrices are denoted by $\mathbb{R}^{N_{1}\times N_{2}}$ and $\mathbb{C}^{N_{1}\times N_{2}}$, respectively. The $l_{2}$-norm of vector $\boldsymbol{v}$ is denoted by $\lVert\boldsymbol{v}\rVert$. For a real number $c$, $\lvert c\rvert$ denotes its absolute value and $\lfloor c\rfloor$ denotes the maximum integer that is no larger than $c$. The statistical expectation is denoted by $\mathbb{E}\{\cdot\}$. $\boldsymbol{I}_{N}$ denotes the $N$-dimensional identity matrix.
@@ II System Model
In the sensing process, we consider that the MA array emits a sensing signal and receives the echoes reflected from the target across T snapshots, during which the target is assumed to remain static [18, 21]. To characterize the near-field channel model from the antenna array to the target, we adopt a uniform spherical wave (USW) channel model in [19], where the path loss of the channel coefficient with respect to (w.r.t.) all MAs is identical while the phase varies over them. As depicted in Fig. 1, we denote the physical steering angle from the origin to the target as $\theta$, while the spatial AoA is defined as $u\triangleq\cos\theta\in[0,1)$ [^1] We exclude the case with $u<0$ as it is equivalent to placing the array in the other side, which yields the same sensing result as the case with $u>0$ due to the symmetry.. Denote $\boldsymbol{s}_{n}=[x_{n},0]^{\top}$ as the coordinate of the $n$-th MA, and $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target, where $r_{\text{min}}$ and $r_{\text{max}}$ are the prescribed lower-bound and upper-bound of the distance range, respectively. Thus, the coordinate of the target is given by $\boldsymbol{r}=[r\cos\theta,r\sin\theta]^{\top}$. Then, the distance from the $n$-th MA to the target can be expressed as a function of the APV $\boldsymbol{x}$ and the target parameters denoted by $\boldsymbol{\eta}=[u,r]^{\top}$, i.e.,
@@ III-A AoA Estimation
where $\tilde{\boldsymbol{x}}\triangleq[\tilde{x}_{1}^{2},\tilde{x}_{2}^{2},\dots,\tilde{x}_{N}^{2}]^{\top}\in\mathbb{R}^{N}$ and $\tilde{x}_{n}\triangleq x_{n}^{2},n\in{\cal N}$. The variance functions are defined as $\text{var}(\boldsymbol{x})\triangleq\frac{1}{N}\sum_{n=1}^{N}x_{n}^{2}-\mu(\boldsymbol{x})^{2}$ with $\mu(\boldsymbol{x})=\frac{1}{N}\sum_{n=1}^{N}x_{n}$ being the mean of $\boldsymbol{x}$, and $\text{var}(\tilde{\boldsymbol{x}})\triangleq\frac{1}{N}\sum_{n=1}^{N}\tilde{x}_{n}^{2}-\mu(\tilde{\boldsymbol{x}})^{2}$ with $\mu(\tilde{\boldsymbol{x}})=\frac{1}{N}\sum_{n=1}^{N}\tilde{x}_{n}$ being the mean of $\tilde{\boldsymbol{x}}$. The covariance function is defined as $\text{cov}(\boldsymbol{x},\tilde{\boldsymbol{x}})\triangleq\frac{1}{N}\sum_{n=1}^{N}x_{n}\tilde{x}_{n}-\mu(\boldsymbol{x})\mu(\tilde{\boldsymbol{x}})$.

### v3

@@ I Introduction
In future sixth-generation (6G) wireless systems, significant advancement in their communication and sensing capabilities is envisioned[1, 2]. A multitude of research efforts have been dedicated to achieving ultra-high transmission rates and precise information acquisition from the physical environment, driven by the demands of emerging applications such as smart healthcare, vehicle-to-everything (V2X) and virtual reality (VR) [3, 4]. However, existing communication and/or sensing systems typically deploy fixed-position antennas (FPAs) at the transmitter/receiver (Tx/Rx), which cannot fully exploit the spatial degrees of freedom (DoFs).
@@
Notations: Boldface lower and upper cases are used to represent vectors and matrices, respectively. The conjugate of a complex is denoted by $(\cdot)^{*}$. The transpose and conjugate transpose of a matrix/vector are denoted by $(\cdot)^{\top}$ and $(\cdot)^{\mathrm{H}}$, respectively. The sets of $N_{1}\times N_{2}$ dimensional real and complex matrices are denoted by $\mathbb{R}^{N_{1}\times N_{2}}$ and $\mathbb{C}^{N_{1}\times N_{2}}$, respectively. The $l_{2}$-norm of vector $\boldsymbol{v}$ is denoted by $\lVert\boldsymbol{v}\rVert$. For a real number $c$, $\lvert c\rvert$ denotes its absolute value and $\lfloor c\rfloor$ denotes the maximum integer that is no larger than $c$. The statistical expectation is denoted by $\mathbb{E}\{\cdot\}$. $\boldsymbol{I}_{N}$ denotes the $N$-dimensional identity matrix.
@@ II System Model
In the sensing process, we consider that the MA array emits a sensing signal and receives the echoes reflected from the target across T snapshots, during which the target is assumed to remain static [18, 21]. To characterize the near-field channel model from the antenna array to the target, we adopt a uniform spherical wave (USW) channel model in [19], where the path loss of the channel coefficient with respect to (w.r.t.) all MAs is identical while the phase varies over them. As depicted in Fig. 1, we denote the physical steering angle from the origin to the target as $\theta$, while the spatial AoA is defined as $u\triangleq\cos\theta\in[0,1)$ [^1] We exclude the case with $u<0$, as it is equivalent to placing the array in the other side, which yields the same sensing result as the case with $u>0$ due to symmetry.. Denote $\boldsymbol{s}_{n}=[x_{n},0]^{\top}$ as the coordinate of the $n$-th MA, and $r\in[r_{\text{min}},r_{\text{max}}]$ as the distance between the origin and the target, where $r_{\text{min}}$ and $r_{\text{max}}$ are the prescribed lower-bound and upper-bound of the distance range, respectively. Thus, the coordinate of the target is given by $\boldsymbol{r}=[r\cos\theta,r\sin\theta]^{\top}$. Then, the distance from the $n$-th MA to the target can be expressed as a function of the APV $\boldsymbol{x}$ and the target parameters denoted by $\boldsymbol{\eta}=[u,r]^{\top}$, i.e.,
@@ III-A AoA Estimation
where $\tilde{\boldsymbol{x}}\triangleq[\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{N}]^{\top}\in\mathbb{R}^{N}$ and $\tilde{x}_{n}\triangleq x_{n}^{2},n\in{\cal N}$. The variance functions are defined as $\text{var}(\boldsymbol{x})\triangleq\frac{1}{N}\sum_{n=1}^{N}x_{n}^{2}-\mu(\boldsymbol{x})^{2}$ with $\mu(\boldsymbol{x})=\frac{1}{N}\sum_{n=1}^{N}x_{n}$ being the mean of $\boldsymbol{x}$, and $\text{var}(\tilde{\boldsymbol{x}})\triangleq\frac{1}{N}\sum_{n=1}^{N}\tilde{x}_{n}^{2}-\mu(\tilde{\boldsymbol{x}})^{2}$ with $\mu(\tilde{\boldsymbol{x}})=\frac{1}{N}\sum_{n=1}^{N}\tilde{x}_{n}$ being the mean of $\tilde{\boldsymbol{x}}$. The covariance function is defined as $\text{cov}(\boldsymbol{x},\tilde{\boldsymbol{x}})\triangleq\frac{1}{N}\sum_{n=1}^{N}x_{n}\tilde{x}_{n}-\mu(\boldsymbol{x})\mu(\tilde{\boldsymbol{x}})$.